write in polar form: negative sqroot of 5+i
Found 2 solutions by fractalier, venugopalramana:
Answer by fractalier(6550)

r = sqrt(x^2 + y^2)
theta = arctan(y/x) ...I use q for theta
so that
r = sqrt(5 + 1) = sqrt(6)
q = arctan(1/-sqrt(5))
and thus in polar form we have
(sqrt(6), arctan(1/-sqrt(5)) or about
(2.449, -24.1°)

Answer by venugopalramana(3286)

write in polar form: negative sqroot of 5+i
Z=5+I
|Z|=R=SQRT(5^2+1^2)=SQRT(26)
TAN(THETA)=1/5.....THETA=TAN^-1 (1/5)=A SAY
Z=R*E^(iA)={(26)^0.5}*{E^(iA)}
-SQRT(Z)=-{(26)^0.25}*E^(iA/2)=-2.258*E^(0.1974i)....ANGLE IN RADIANS ..OR...
=-2.258*E^(11.316i)....ANGLE IN DEGREES